With the help of triple integrals, find the volume of the sphere ρ = 2 in a) Rectangular coordianates b) Cylinderical coordiantes c) Spherical coordinates 



1
Expert's answer
2021-03-31T07:32:05-0400

With the help of triple integrals, find the volume of the sphere ρ = 2 in a) Rectangular coordianates b) Cylinderical coordiantes c) Spherical coordinates 

Solution:

a) Rectangular coordinates:


"x^2+y^2+z^2=4"

"V=\displaystyle\int_{-2}^2dx\int_{-\sqrt{4-x^2}}^{\sqrt{4-x^2}}dy\int_{-\sqrt{4-x^2-y^2}}^{\sqrt{4-x^2-y^2}}dz=" "8\displaystyle\int_{0}^2dx\int_{0}^{\sqrt{4-x^2}}dy\int_{0}^{\sqrt{4-x^2-y^2}}dz=" "8\displaystyle\int_{0}^2dx\int_{0}^{\sqrt{4-x^2}}\sqrt{4-x^2-y^2}dy"

Let's apply the change of variables:

"y=\sqrt{4-x^2}\cdot\sin{t}"

"dy=\sqrt{4-x^2}\cdot\cos{t}\cdot dt"

"y=0 \implies t=0"

"y=\sqrt{4-x^2}\implies t=\frac{\pi}{2}"

"V=8\displaystyle\int_{0}^2dx\int_{0}^{\pi/2}\sqrt{4-x^2-(4-x^2)\sin^2{t}}\cdot" "\sqrt{4-x^2}\cos{t}dt="

"8\displaystyle\int_{0}^2dx\int_{0}^{\pi/2}(4-x^2)\cos^2{t}dt=" "8\displaystyle\int_{0}^2dx\int_{0}^{\pi/2}(4-x^2)\frac{(1+\cos{2t})}{2}dt=" "8\displaystyle\int_{0}^2dx(4-x^2)\frac{(t+\frac{\sin{2t}}{2})}{2}|_0^{\pi/2}=" "2\pi\displaystyle\int_{0}^2(4-x^2)dx=2\pi(4x-x^3/3)|_0^2=\frac{32\pi}{3}"

b) Cylinderical coordiantes:

"x=\rho\cos{\varphi}"

"y=\rho\sin{\varphi}"

"z=z"

"J=\rho"

The equation of the sphere:

"\rho^2+z^2=4"

"V=\displaystyle\int_0^{2\pi}d\varphi\int_0^2d\rho\int_{-\sqrt{4-\rho^2}}^{\sqrt{4-\rho^2}}\rho dz=" "2\displaystyle\int_0^{2\pi}d\varphi\int_0^2d\rho\int_0^{\sqrt{4-\rho^2}}\rho dz=" "4\pi\displaystyle\int_0^2\sqrt{4-\rho^2}\rho d\rho=" "2\pi\displaystyle\int_0^2\sqrt{4-\rho^2} d\rho^2=-\frac{4\pi}{3}\displaystyle(4-\rho^2)^{3/2}|_0^2=\frac{32\pi}{3}"

c) Spherical coordinates 

"x=r\sin{\theta}\cos{\varphi}"

"y=r\sin{\theta}\sin{\varphi}"

"z=r\cos{\theta}"

"J=r^2\sin{\theta}"

"V=\displaystyle\int_0^{2\pi}d\varphi\int_0^{\pi}d\theta\int_0^2r^2\sin{\theta}dr=" "2\pi\displaystyle\int_0^{\pi}d\theta\sin{\theta}\frac{r^3}{3}|_0^2=" "\displaystyle\frac{16\pi}{3}\int_0^{\pi}\sin{\theta}d\theta=" "-\displaystyle\frac{16\pi}{3}\cos{\theta}|_0^{\pi}=\frac{32\pi}{3}"

Answer: "V=\displaystyle\frac{32\pi}{3}."